Air within the funnel of a large tornado may
have a pressure of only 0.19 atm.
What is the approximate outward force F
on a 9.1 m × 26 m wall if a tornado suddenly
envelopes the house? Atmospheric pressure is 1.013 × 105 Pa.
Answer in units of N.
1
Expert's answer
2011-01-25T09:27:31-0500
As in 1 atm is 1.013 × 105 Pa, in 0.19 atm would be 0.192 × 105 Pa.
Let's find the area of the wall: S = 9.1* 26 = 236.6 m2.
Thus:
F = PS = 0.192 × 105 * 236.6 = 4.543 × 106 N.